# Simulating the Real World: A Unified Survey of Multimodal Generative Models

### v1

@@ 3.3.2 Algorithms
TABLE II: Recent text-to-3D, image-to-3D and video-to-3D generation methods.
@@
[…]
@@
[…]
@@
(1) Text-to-3D Generation. To generate 3D content from the text prompt by simulating the geometrics in the real world, extensive studies have been conducted and can be divided into three branch. Readers can refer to [161, 162, 163] for a more comprehensive survey in this field.
@@
(2) Image-to-3D Generation. The goal of image-to-3D task is to generate high-quality 3D assets that are consistent with the identity of the given image.
(i) Feedforward Approaches. These methods first encode 3D assets into latent codes with a compression network, like VAE, and then train a generative model for the latent space samples.
3DGen[139] introduce triplane as latent space, improving both the accuracy and efficiency of the compression network. Direct3D[142] adopt the triplane representation and directly use 3D supervision for training, preserving detailed 3D information in the latent triplane. Michelangelo[122] draws inspiration from 3Dshape2vecset[166], and utilize 1D vectors as the latent space, when supervise the output with occupancy fields. CraftsMan[141] further introduces multi-view generation model to generate multi-view images as conditions of the diffusion model, followed with normal-based refinement for generated meshes. Clay[140] brings a comprehensive system pretrained on large-scale 3D datasets for 3D generation, including VAE and diffusion model based on 1D vectors for geometry generation, Material Diffusion for PBR texture, and conditioning design across various modalities.
(ii) Optimization-based Approaches. With the development of distillation-based methods in text-3D models, optimization-based methods directly optimize 3D assets through a training process supervised by SDS loss from pretrained image-image or text-image generation models, when maintaining the image identity by various additional loss constraints.
Adapted from Dreamfusion[127], Magic3D[124] and SJC [171], RealFusion[144] only extracts priors from pretrained text-image models with the SDS loss, when the image reconstruction loss and textual inversion is used to preserve low-level and semantic identity respectively. With the advent of large-scale open-set 3D datasets[167], Zero123[145] introduces the priors of 3D datasets into image-3D tasks by replacing the text-image model with a novel view synthesis model for distillation. Specifically, Zero123 finetunes a pretrained image-image generation model on the 3D datasets by introducing camera poses as a condition to control the generated image’s viewpoint. The pretrained novel view synthesis model preserves both the details from the image models and the Multi-view consistency from 3D datasets, which significantly mitigates the Janus problem.
A series of works have expanded upon Zero123[145]. Zero123-xl[168] pretrains Zero123 pipelines on a 10$\times$ larger 3D datasets for better generalization. Magic123[146] leverages 2D and 3D priors simultaneously for distillation to manage the trade-off between generalization and consistency, and uses coarse-fine pipeline for higher quality. SyncDreamer[147] and Consistent123[148] both further improves the multi-view consistency of NVS models by introducing a synchronized multi-view diffusion model, when the former leverages 3D volume to model joint distribution relations of images and the latter utilizes cross-view attention and shared self-attention. Toss[149] additionally brings text caption as high-level semantics of 3D data into th NVS model pretaining for stronger plausibility and controllability of invisible views. ImageDream[150] addresses both multi-view consistency and 3D details problem by designing a multi-level image-prompt controller and training with text descriptions. Wonder3D[151] incorporates the cross-domain attention mechanism, enabling the NVS model to simultaneously denoise images and align normal maps, when introducing normal maps into the optimization process additionally.
@@
(iii) MVS-based Approaches. MVS-based methods split the image-3D generation into two stages: first generate multi-view images from a single image using an NVS model, then directly create 3D assets from these multi-view images using a feedforward reconstruction network.
Based on the multi-view images predicted by Zero123[145], One-2-3-45[152] proposes an elevation estimation module and utilizes an SDF-based generalizable neural surface reconstruction module pretrained on 3D datasets for 360◦ mesh reconstruction, which reduces the reconstruction time to 45 seconds compared to optimization-based. CRM[153] further freeze the output images of the multi-view generation model to six fixed camera poses, largely improving the consistency between multi-views. Then CRM input multi-view images into a convolutional U-Net to create a high-resolution triplane supervised by depth and RGB images. InstantMesh[154] also freeze the camera poses of multi-view images, but employ a transformer-based multi-view reconstruction model based on LRM[155] to reconstruct 3D meshes, providing better generalization at the expense of some image-3D detail consistency. Unique3d[156] introduce a multi-level upscale strategy to progressively generate higher-resolution multi-view images, and a normal map diffusion model to predict multi-view normal maps for the initialization of coarse meshes, which is refined and colorized based on multi-view images.
@@
(3) Video-to-3D Generation. The vast corpus of online video data serves as a rich repository of 3D information, capturing object motions, viewpoint variations, and camera transitions that unveil multi-view perspectives often unavailable in static imagery [172, 173, 174, 175, 176, 177]. This dynamic content provides both temporal coherence and spatial consistency across sequential frames, which are essential for understanding complex 3D scenes and generating high-fidelity 3D structures [13]. As a result, leveraging these multi-view and time-variant data has become a promising approach for reconstructing and synthesizing 3D-consistent objects [178]. Recent research explores video-based priors for robust 3D generation  [160, 158, 159], aiming to learn 3D representations that maintain coherence across frames and adapt to changing viewpoints. At high-level, the main idea of these works for video-to-3D generation is to enable a camera-controllable video model as a consistent multi-view generator for dense 3D reconstruction.
@@ 3.3.3 Applications
3.4 4D Generation
We culminate in 4D generation by integrating all dimensions. As a cutting-edge field in computer vision, 4D generation focuses on synthesizing dynamic 3D scenes that evolve temporally based on multimodal inputs such as text, images, or videos. Unlike traditional 2D or 3D generation [208], 4D synthesis introduces unique challenges, requiring both spatial coherence and temporal consistency while balancing high fidelity, computational efficiency, and dynamic realism [209]. In this section, we first introduce 4D representation, which builds upon 3D representation, and then summarize current 4D generation methods. Recent research has explored two primary paradigms: optimization-based methods leveraging Score Distillation Sampling (SDS) and feedforward-based approaches that avoid per-prompt optimization. These paradigms address distinct technical challenges, highlighting the field’s complexity and the ongoing pursuit of a feasible balance among visual quality, computational efficiency, and scene flexibility. Representative works in 4D generation are summarized in Table III.
TABLE III: Representative works of 4D generation methods. ”Rep” stands for representations.
@@
[…]
@@
3.4.1 4D Representation
The field of 4D representation, which incorporates a temporal dimension into 3D modeling, provides a strong foundation for understanding dynamic scenes. By extending static 3D spatial representations $(x,y,z)$ with time $(t)$, these methods encode scene dynamics and transformations, essential for applications such as non-rigid human motion capture and simulating object trajectories [230, 231, 232, 233].
@@
3.4.2 Algorithms
(1) Feedforward Approaches. Feedforward-based methods provide an efficient alternative by generating 4D content in a single forward pass, bypassing the iterative optimization required in SDS-based pipelines. These methods rely on pre-trained models, leveraging temporal and spatial priors to achieve fast and consistent generation. Control4D [210] and Animate3D [211] directly synthesize dynamic scenes from textual or visual inputs, enabling real-time applications such as interactive media and personalized content creation. Vidu4D [212] improves motion trajectories by incorporating temporal priors, ensuring frame-to-frame coherence and smooth transitions. Diffusion4D [213] extends the capabilities of diffusion models to handle 4D scene synthesis by combining spatial-temporal feature extraction with efficient inference mechanisms. L4GM [214] further enhances feedforward techniques by integrating latent geometry modeling, producing high-quality results while maintaining computational efficiency.
@@ 3.4.2 Algorithms
(2) Optimization-based Approaches. Optimization-based methods are foundational in 4D generation, employing iterative techniques such as Score Distillation Sampling (SDS) to adapt pre-trained diffusion models to synthesize dynamic 4D scenes. These approaches leverage powerful priors from text-to-image, multi-view image, and text-to-video generation models, achieving temporally coherent scenes with rich motion dynamics. For example, MAV3D [215] optimizes NeRF or HexPlane features against SDS loss guided by textual prompts, while 4D-fy [216] and Dream-in-4D [219] improve the 3D consistency and motion dynamics by integrating image, multi-view, and video diffusion models in SDS supervision. AYG [217] proposes to use deformable 3DGS as inherent representation, which can easily disentangle static geometry from dynamic motion with a simple delta deformation field to improve flexibility.
@@
TABLE IV: Summary of the widely-used 2D, video, 3D and 4D generation datasets. [Link] directs to dataset websites.
@@
Dataset Year Numbers Type Evaluation Metrics Contribution Highlights
2D Generation (Sec. 3.1)
@@
SBU [242] [Link] 2011 1M Image-Text pairs - Photographs with captions and 5 kinds of image content.
@@
MS-COCO [243] [Link] 2014 330K Image-Text pairs FID, CLIP-Score, SSIM Manual annotation and 330K for English.
@@
CC-3M [244] [Link] 2018 3M Image-Text pairs - Alt-text caption annotation.
@@
LAION-5B [169] [Link] 2022 5.85B Image-Text pairs FID, CLIP-Score, SSIM 5.85B CLIP-filtered multilingual image-text pairs.
@@
ShareGPT4V  [245] [Link] 2023 102K Image-Text pairs - 1.2M image-text pairs for ShareGPT4V-PT.
@@
Pick-a-Pic [246] [Link] 2023 500K(35K distinct prompts) Image-Text pairs FID, PickScore Human preferences on model-generated images.
Video Generation (Sec. 3.2)
@@
UCF-101 [247] [Link] 2012 13K action recognition FVD Manual caption annotation.
@@
ActivityNet [248] [Link] 2015 28K action recognition FVD Manual caption annotation.
@@
MSR-VTT [249] [Link] 2016 10K video-language dataset FVD, CLIP-Score Manual caption annotation. Resolution: 240p.
@@
HowTo100M [250] [Link] 2019 136M video-language dataset FVD, CLIP-Score ASR annotation. Resolution: 240p.
@@
WebVideo-10M [251] 2021 10M video-language dataset FVD, CLIP-Score Alt-text caption annotation. Resolution: 360p.
@@
HD-VILA-100M [252] [Link] 2022 103M video-language dataset FVD, CLIP-Score ASR annotation. Resolution: 720p.
@@
InternVid [253] [Link] 2023 7M video-language dataset FVD, CLIP-Score Automatic caption annotation.
@@
Panda-70M [170] [Link] 2024 70M video-language dataset FVD, CLIP-Score Automatic caption annotation. Resolution: 720p.
@@
Koala-36M [254] [Link] 2024 36M video-language dataset FVD, CLIP-Score Automatic and manual caption annotation. Resolution: 720p.
3D Generation (Sec. 3.3)
@@
DeepFashion [255] [Link] 2016 800K single-view images - Human clothing dataset.
@@
SHHQ [256] [Link] 2022 40K single-view images - Human body dataset.
CO3D [257] [Link] 2021 19K(1.5M images) multi-view images PSNR/LPIPS; IoU, etc. Image-to-3D Generation. 50 classes. Annotations: cameras, pcl.
RTMV [258] [Link] 2021 300K multi-view images PSNR/LPIPS; MAE, etc. 2,000 complex scenes with ray tracing (1600×1600).
MVImgNet [259] [Link] 2023 219K multi-view images PSNR/LPIPS 6.5M frames, 238 classes. Annotations: cameras, pcl.
uCO3D [260] [Link] 2025 170K multi-view images PSNR/LPIPS; IoU 1K classes. Annotations: cameras, 3DGS, captions.
ShapeNet [261] [Link] 2015 51K 3D data - 51K 3D CAD models across 3K object categories.
@@
3D-Future [262] [Link] 2020 10K 3D data CLIP Similarity 9,992 3D furniture models with textures and attributes.
@@
GSO [263] [Link] 2022 1K 3D data F-Score/ CLIP Similarity Image-to-3D Generation. 1,030 scanned household objects.
@@
Objaverse [167] [Link] 2022 800K 3D data F-Score/ CLIP Similarity Objaverse and Objaverse-1.0 are the same.
@@
Objaverse-XL [168] [Link] 2023 10.2M 3D data F-Score/ CLIP Similarity Image-to-3D Generation.
Cap3D [264] [Link] 2023 785k 3D-Text pairs FID/CLIP etc. Text-to-3D Generation. Auto/manual captions.
@@
Animal2400 [116] 2023 Training(300), Testing(2400) prompts CLIP R-Precision Text-to-3D Generation. 10 animals, 8 activities, 6 themes, 5 hats.
DreamFusion [127] [Link] 2024 DF27/DF415 prompts CLIP R-Precision Text-to-3D Generation. DF27: 27 paper prompts; DF415: 415 webpage prompts.
4D Generation (Sec. 3.4)
@@
Consistent4D [223] [Link] 2023 12(in-the-wild), 14(synthetic) multi-view videos LPIPS/CLIP Video-to-4D Generation.
Diffusion4D [213] [Link] 2024 365K dynamic 3D data - Collected from Objaverse-1.0 (42K) and Objaverse-xl (323K)
@@
MV-Video [211] [Link] 2024 53K dynamic 3D data - Annotated by MiniGPT4-Video [265]
CamVid-30K [266] [Link] 2024 30K 4D data - Real-world 4D scene dataset
@@
4D-DRESS [267] [Link] 2024 64(outfits) human clothing 4D data CD Total number of 3D human frames: 78k.
@@
TABLE V: Summary of common evaluation metrics.
@@
[…]
3.4.3 Applications
(2) Human Animation. As an essential component of 4D simulation, human motion generation is a research branch that attracts the most attention from the community. Different from the human-centric video generation (in section 3.2.2), 3D human motion generation is easier to animate characters in 3D applications, e.g. games and embodied intelligence. The recent success of the 3D human motion generation mainly comes from the well-studied human body parameterized models [283, 284].
4 Datasets and Evaluations

### v2

@@ III-C2 Algorithms
TABLE IV: Recent text-to-3D, image-to-3D and video-to-3D generation methods.
@@
Optimization Magic3D [144] Link Link NeRF/DMTet SDS
@@
Optimization RealFusion [165] Link Link NeRF SDS/Image Reconstruction
@@
(2) Image-to-3D Generation. The goal of image-to-3D task is to generate high-quality 3D assets that are consistent with the identity of the given image. Due to the high cost of 3D data collection, text-to-3D generation lacks sufficient high-quality text annotations to scale up compared with image and video generation. Since images naturally capture more low-level information that closely align with the 3D modality, the image-to-3D task narrows the modality gap between the input and the output compared to text-to-3D generation. Consequently, image-to-3D has emerged as a foundational task for advancing native 3D generation. To leverage the knowledge inside image generation models, they are frequently adopted as components of image-to-3D models (Fig. 8). The qualitative comparison of the part methods is shown in Fig. 9 and the quantitative comparison is illustrated in Table V. Due to inconsistencies in the evaluation datasets or metrics used in the paper, some works are not listed in the figure and table.
Figure 8: Comparison between different image-to-3D generation paradigms. Image generation models can be employed to synthesize multi-view images to promote 3D generation.
Figure 9: Qualitative comparison between different image-to-3D methods.
TABLE V: Quantitative comparison with baseline methods on the GSO dataset [183].
@@
Optimization-based Approaches Realfusion [165] 0.0819 0.2741 ~90min
Magic123 [167] 0.0516 0.4528 ~60min
MVS-based Approaches One-2-3-45 [174] 0.0629 0.4086 ~45s
Zero123 [166] 0.0339 0.5035 ~10min
InstantMesh [176] 0.0187 0.6353 ~10s
Feedforward Approaches Point-E [135] 0.0426 0.2875 ~40s
Shap-E [138] 0.0436 0.3584 ~10s
Michelangelo [131] 0.0404 0.4002 ~3s
CraftsMan3D[162] 0.0291 0.5347 ~5s
(i) Feedforward Approaches. These methods first encode 3D assets into latent codes with a compression network, like VAE, and then train a generative model for the latent space samples. 3DGen[160] introduce triplane as latent space, improving both the accuracy and e fficiency of the compression network. Direct3D[163] adopt the triplane representation and directly use 3D supervision for training, preserving detailed 3D information in the latent triplane. Michelangelo[131] draws inspiration from 3Dshape2vecset[184], and utilize 1D vectors as the latent space, when supervise the output with occupancy fields. CraftsMan[162] further introduces multi-view generation model to generate multi-view images as conditions of the diffusion model, followed with normal-based refinement for generated meshes. Clay[161] brings a comprehensive system pretrained on large-scale 3D datasets for 3D generation, including VAE and diffusion model based on 1D vectors for geometry generation, Material Diffusion for PBR texture, and conditioning design across various modalities.
(ii) Optimization-based Approaches. With the development of distillation-based methods in text-to-3D models, optimization-based methods directly optimize 3D assets through a training process supervised by SDS loss from pretrained image-to-image or text-to-image generation models, when maintaining the image identity by various additional loss constraints.
Adapted from Dreamfusion[142], Magic3D[144] and SJC [189], RealFusion[165] only extracts priors from pretrained text-to-image models with the SDS loss, when the image reconstruction loss and textual inversion is used to preserve low-level and semantic identity respectively. With the advent of large-scale open-set 3D datasets[185], Zero123[166] introduces the priors of 3D datasets into image-to-3D tasks by replacing the text-to-image model with a novel view synthesis model for distillation. Specifically, Zero123 finetunes a pretrained image-image generation model on the 3D datasets by introducing camera poses as a condition to control the generated image’s viewpoint. The pretrained novel view synthesis model preserves both the details from the image models and the multi-view consistency from 3D datasets, which significantly mitigates the Janus problem.
@@
(iii) MVS-based Approaches. MVS-based methods split the image-to-3D generation into two stages: first generate multi-view images from a single image using an NVS model, then directly create 3D assets from these multi-view images using a feedforward reconstruction network.
Based on the multi-view images predicted by Zero123[166], One-2-3-45[174] proposes an elevation estimation module and utilizes an SDF-based generalizable neural surface reconstruction module pretrained on 3D datasets for $360^{\circ}$ mesh reconstruction, which reduces the reconstruction time to 45 seconds compared to optimization-based. CRM[175] further freeze the output images of the multi-view generation model to six fixed camera poses, largely improving the consistency between multi-views. Then CRM input multi-view images into a convolutional U-Net to create a high-resolution triplane supervised by depth and RGB images. InstantMesh[176] also freeze the camera poses of multi-view images, but employ a transformer-based multi-view reconstruction model based on LRM[177] to reconstruct 3D meshes, providing better generalization at the expense of some image-to-3D detail consistency. Unique3d[178] introduce a multi-level upscale strategy to progressively generate higher-resolution multi-view images, and a normal map diffusion model to predict multi-view normal maps for the initialization of coarse meshes, which is refined and colorized based on multi-view images.
@@
(3) Video-to-3D Generation. Video-to-3D methods fundamentally build upon advances in 2D diffusion models, extending image-based generative priors into the temporal domain. By modeling coherent sequences of frames—with consistent texture, lighting, and geometry—video diffusion enables implicit 3D structure learning through motion and viewpoint variation captured in 2D. Frameworks like SV3D, Hi3D, and V3D adapt pretrained video diffusion backbones to generate multi‑view frame sequences, which are then integrated via volume rendering, mesh optimization or Gaussian splatting to reconstruct explicit 3D geometry. This synergy leverages the strong content synthesis capabilities of 2D video models while introducing spatial consistency and camera control essential for dense 3D reconstruction. The vast corpus of online video data serves as a rich repository of 3D information, capturing object motions, viewpoint variations, and camera transitions that unveil multi-view perspectives often unavailable in static imagery [190, 191, 192, 193, 194, 195]. This dynamic content provides both temporal coherence and spatial consistency across sequential frames, which are essential for understanding complex 3D scenes and generating high-fidelity 3D structures [13]. As a result, leveraging these multi-view and time-variant data has become a promising approach for reconstructing and synthesizing 3D-consistent objects [196]. Recent research explores video-based priors for robust 3D generation  [182, 180, 181], aiming to learn 3D representations that maintain coherence across frames and adapt to changing viewpoints. At high-level, the main idea of these works for video-to-3D generation is to enable a camera-controllable video model as a consistent multi-view generator for dense 3D reconstruction (Fig. 10).
Figure 10: An illustration of video-to-3D generation paradigm. Camera-controllable video models can be used as a consistent multi-view generator for 3D generation.
@@ III-C3 Applications
III-D 4D Generation
We culminate in 4D generation by integrating all dimensions. As a cutting-edge field in computer vision, 4D generation focuses on synthesizing dynamic 3D scenes that evolve temporally based on multimodal inputs such as text, images, or videos. Unlike traditional 2D or 3D generation [226], 4D synthesis introduces unique challenges, requiring both spatial coherence and temporal consistency while balancing high fidelity, computational efficiency, and dynamic realism [227]. In this section, we first introduce 4D representation, which builds upon 3D representation, and then summarize current 4D generation methods. Recent research has explored two primary paradigms: optimization-based methods leveraging Score Distillation Sampling (SDS) and feedforward-based approaches that avoid per-prompt optimization. These paradigms address distinct technical challenges, highlighting the field’s complexity and the ongoing pursuit of a feasible balance among visual quality, computational efficiency, and scene flexibility. Representative works in 4D generation are summarized in Table VI.
TABLE VI: Representative works of 4D generation methods. ”Rep” stands for representations.
@@
GenXD [233] Link Link 4DGS Diffusion-based SSIM+LPIPS
CAT4D [234] Link - Deformable 3DGS Diffusion-based LPIPS
4Real-Video [235] Link - Deformable 3DGS Diffusion-based velocity matching loss of rectified flow [236]
Optimization MAV3D [237] Link - HexPlane-/NeRF-based 4D Rep T2I/T2V SDS
@@
4K4DGen [252] Link Link 4DGS SVD L1
EG4D [253] Link Link 4DGS SVD L1
AvatarGO [254] Link Link 3DGS+Hexplane SVD/Zero123 SDS
Disco4D [255] Link Link 4DGS SVD SDS
GenMOJO [256] Link Link Deformable 3DGS SVD/Zero123 SDS
III-D1 4D Representation
The field of 4D representation, which incorporates a temporal dimension into 3D modeling, provides a strong foundation for understanding dynamic scenes. By extending static 3D spatial representations $(x,y,z)$ with time $(t)$, these methods encode scene dynamics and transformations, essential for applications such as non-rigid human motion capture and simulating object trajectories [257, 258, 259, 260]. Most 4D representations can be factorized into canonical 3d representation and deformation. The first module canonical 3d representation is used to model a static template shape. The second module deformation is used to deform the template shape to synthesize motion. To deform the canonical 3d representation, here are some common used deformation representations including deformation fields that is a neural networks that map a space-time point to its corresponding position on the canonical template shape and deformation primitives like linear blend skinning which represents the motion of a point through a combination of rigid motions associated with different body parts or control points. There are pros and cons to these two types of deformation representations. Deformation fields are more flexible than deformation primitives, which means deformation fields are more general and can theoretically fit complex motions. Deformation primitives designed for articulated objects such as human or animal motions, however, have the advantage of robustness for large articulated motion since deformation fields lack the inductive bias needed for accurate reconstruction, especially with rapid motion. Below, we focus most on the 3d canonical representation with deformation fields type of 4D representation.
@@
III-D2 Algorithms
Modern 4D generation methods are largely rooted in the foundations laid by 3D generation. In particular, advances in 3D reconstruction, such as NeRF and 3DGS, have directly influenced how we model and render dynamic 4D scenes. These 3D frameworks provide not only efficient data structures and rendering techniques but also critical inductive biases that benefit temporal modeling in 4D. At the representation level, 3D methods offer canonical spatial priors that can be extended with deformation fields or motion trajectories to capture temporal evolution. From a training perspective, fast-training techniques from 3D (e.g., hash encoding, hierarchical sampling) have been adapted to accelerate 4D optimization. Human animation, as a most representative 4D task, particularly benefits from foundations and advancements in 3D human modeling. Techniques such as SMPL(-X), linear blend skinning, and neural deformation fields provide strong structural priors for modeling articulated motion. These 3D-based tools enhance both the realism and controllability of 4D human motion synthesis.
Figure 11: An illustration of 4D generation feedforward paradigm. We directly train the 4D generator on the multi-view dynamic data.
(1) Feedforward Approaches. Feedforward-based methods provide an efficient alternative by generating 4D content in a single forward pass (Fig. 11), bypassing the iterative optimization required in SDS-based pipelines. These methods rely on pre-trained models, leveraging temporal and spatial priors to achieve fast and consistent generation. Control4D [228] and Animate3D [229] directly synthesize dynamic scenes from textual or visual inputs, enabling real-time applications such as interactive media and personalized content creation. Vidu4D [230] improves motion trajectories by incorporating temporal priors, ensuring frame-to-frame coherence and smooth transitions. Diffusion4D [231] extends the capabilities of diffusion models to handle 4D scene synthesis by combining spatial-temporal feature extraction with efficient inference mechanisms. L4GM [232] further enhances feedforward techniques by integrating latent geometry modeling, producing high-quality results while maintaining computational efficiency.
@@ III-D2 Algorithms
Figure 12: An illustration of 4D generation optimization-based paradigm. We use SDS guidance to make the optimized 4D asset follow the pretrained text-to-image, multi-view image, and text-to-video prior.
(2) Optimization-based Approaches. Optimization-based methods are foundational in 4D generation, employing iterative techniques such as Score Distillation Sampling (SDS) to adapt pre-trained diffusion models to synthesize dynamic 4D scenes (Fig. 12). These approaches leverage powerful priors from text-to-image, multi-view image, and text-to-video generation models, achieving temporally coherent scenes with rich motion dynamics. For example, MAV3D [237] optimizes NeRF or HexPlane features against SDS loss guided by textual prompts, while 4D-fy [238] and Dream-in-4D [241] improve the 3D consistency and motion dynamics by integrating image, multi-view, and video diffusion models in SDS supervision. AYG [239] proposes to use deformable 3DGS as inherent representation, which can easily disentangle static geometry from dynamic motion with a simple delta deformation field to improve flexibility.
@@
Evaluation. Quantitative evaluation of 4D generation methods mainly focuses on three aspects: 1) Time cost to generate a single 4D asset. 2) CLIP score, which evaluates on image similarity between the generated and the ground truth novel views. 3) Perceptual similarity (LPIPS) between the generated and the ground truth novel views. Here, we report the quantitative scores of representative methods in Table VII and present their visual results in Fig. 13. Readers can refer to [232] for more details.
Time ($\downarrow$) CLIP ($\uparrow$) LPIPS ($\downarrow$)
Consistent4D [245] 2hr 0.87 0.16
4DGen [269] 1hr 0.89 0.13
STAG4D [248] 1hr 0.91 0.13
DG4D [270] 10min 0.87 0.16
TABLE VII: Quantitative comparison between different 4D generation methods. Results are obtained from [232].
Figure 13: Qualitative comparison between different 4D generation methods. Results are obtained from L4GM [232].
III-D3 Applications
(2) Human Animation. As an essential component of 4D simulation, human motion generation is a research branch that attracts the most attention from the community. Different from the human-centric video generation (in Sec. III-B2), 3D human motion generation is easier to animate characters in 3D applications, e.g. games and embodied intelligence. The recent success of the 3D human motion generation mainly comes from the well-studied human body parameterized models [273, 274].
IV Datasets and Evaluations
@@
TABLE VIII: Summary of the widely-used 2D, video, 3D and 4D generation datasets. [Link] directs to dataset websites.
@@
Dataset Year # Categories # Objects # Images Type Evaluation Metrics Contribution Highlights Limitations
2D Generation (Sec. III-A)
@@
SBU [308] [Link] 2011 89 - 1M Image-Text pairs - Photographs with captions and 5 kinds of image content. Noisy user captions and single reference.
@@
MS-COCO [309] [Link] 2014 80 for objects, 91 for stuffs 1.5M 330K Image-Text pairs FID, CLIP-Score, SSIM Manual annotation and 330K for English. Class imbalance and social bias.
@@
CC-3M [310] [Link] 2018 - - 3M Image-Text pairs - Alt-text caption annotation. Noisy user captions and weakly aligned.
@@
LAION-5B [187] [Link] 2022 - - 5.85B Image-Text pairs FID, CLIP-Score, SSIM 5.85B CLIP-filtered multilingual image-text pairs. Uncurated and unsafe contents.
@@
ShareGPT4V  [311] [Link] 2023 - - 102K Image-Text pairs - 1.2M image-text pairs for ShareGPT4V-PT. AI generated captions with hallucination.
@@
Pick-a-Pic [312] [Link] 2023 - - 500K Image-Text pairs FID, PickScore Human preferences on model-generated images. Limited preference labels.
Video Generation (Sec. III-B)
@@
UCF-101 [313] [Link] 2012 101 - 13K action recognition FVD Manual caption annotation. Strong background bias misleads action recognition.
@@
ActivityNet [314] [Link] 2015 200 - 28K action recognition FVD Manual caption annotation. Background bias.
@@
MSR-VTT [315] [Link] 2016 20 - 10K video-language dataset FVD, CLIP-Score Manual caption annotation. Resolution: 240p. Noisy and duplicate captions.
@@
HowTo100M [316] [Link] 2019 12 154 136M video-language dataset FVD, CLIP-Score ASR annotation. Resolution: 240p. Noisy and misaligned narration.
@@
WebVideo-10M [317]  2021 - - 10M video-language dataset FVD, CLIP-Score Alt-text caption annotation. Resolution: 360p. Low-resolution and watermarked videos.
@@
HD-VILA-100M [318] [Link] 2022 15 - 103M video-language dataset FVD, CLIP-Score ASR annotation. Resolution: 720p. Noisy and misaligned ASR subtitles.
@@
InternVid [319] [Link] 2023 16 - 7M video-language dataset FVD, CLIP-Score Automatic caption annotation. Limited diversity.
@@
Panda-70M [188] [Link] 2024 - - 70M video-language dataset FVD, CLIP-Score Automatic caption annotation. Resolution: 720p. Limited diversity.
@@
Koala-36M [320] [Link] 2024 - - 36M video-language dataset FVD, CLIP-Score Automatic and manual caption annotation. Resolution: 720p. Inherent biases from its source corpus.
3D Generation (Sec. III-C)
@@
DeepFashion [321] [Link] 2016 50 300K 800K single-view images - Human clothing dataset. No 3D data.
@@
SHHQ [322] [Link] 2022 - - 230K single-view images - Human body dataset. (1) No 3D data. (2) Inherent biases from FFHQ.
CO3D [323] [Link] 2021 50 19K 1.5M multi-view images PSNR/LPIPS; IoU, etc. Annotated with camera poses and point clouds. Lack high-fidelity geometry/textures.
RTMV [324] [Link] 2021 - 2K 300K multi-view images PSNR/LPIPS; MAE, etc. Large-scale synthetic dataset with high-resolution images. Synthetic data cannot fully capture real-world complexity.
MVImgNet [325] [Link] 2023 238 219K 6.5M multi-view images PSNR/LPIPS Annotated with camera poses and point clouds. Lack high-fidelity geometry/textures.
uCO3D [326] [Link] 2025 1K 170K 170K multi-view images PSNR/LPIPS; IoU Annotated with camera poses, point clouds, depth maps, caption, and 3DGS. Lack high-fidelity geometry/textures.
ShapeNet [327] [Link] 2015 3K 51K 3D data - Large-scale synthetic CAD dataset. Synthetic data cannot fully capture real-world complexity.
@@
3D-Future [328] [Link] 2020 34 10K - 3D data CLIP Similarity High-quality 3D models with textures and attributes. (1) Lack object diversity. (2) Synthetic data cannot fully capture real-world complexity.
@@
GSO [183] [Link] 2022 17 1K - 3D data F-Score/ CLIP Similarity High-fidelity geometry/textures Limited to household objects.
@@
Objaverse [185] [Link] 2022 21K 818K - 3D data F-Score/ CLIP Similarity Large-scale dataset. Variable data quality.
@@
Objaverse-XL [186] [Link] 2023 - 10.2M - 3D data F-Score/ CLIP Similarity Large scale and high diversity. Variable data quality.
Cap3D [329] [Link] 2023 - 785k - 3D-text pairs FID/CLIP etc. Rich text annotations Inherent bias in captions from pretrained models.
@@
Animal2400 [132] 2023 - 2400 - prompts CLIP R-Precision 10 animals, 8 activities, 6 themes, 5 hats. Limited object diversity.
DF27/DF415 [142] [Link] 2024 - 27/415 - prompts CLIP R-Precision - Limited to synthetic or artist-created objects.
4D Generation (Sec. III-D)
@@
Consistent4D [245] [Link] 2023 - - 12(in-the-wild), 14(synthetic) multi-view videos LPIPS/CLIP Video-to-4D Generation. -
Diffusion4D [231] [Link] 2024 - 365K 365K dynamic 3D data - Collected from Objaverse-1.0 (42K) and Objaverse-xl (323K) -
@@
MV-Video [229] [Link] 2024 - 53K 53K dynamic 3D data - Annotated by MiniGPT4-Video [330] Inaccurate captions.
CamVid-30K [233] [Link] 2024 - - 30K 4D data - Real-world 4D scene dataset -
@@
4D-DRESS [331] [Link] 2024 - - 64(outfits) human clothing 4D data CD Total number of 3D human frames: 78k. Manual-heavy pipeline limits scalability.
@@
TABLE IX: Summary of common evaluation metrics.
@@
AQ $\uparrow$ Human Preference. Appearance Quality: percentage unit ($\%$).
GQ $\uparrow$ Human Preference. Geometry Quality: percentage unit ($\%$).
DQ $\uparrow$ Human Preference. Dynamics Quality: percentage unit ($\%$).
TA $\uparrow$ Human Preference. Text Alignment: percentage unit ($\%$).